# Binding and dimerization control phase separation in a compartment

### v1

@@ Abstract
Biological cells exhibit a hierarchical spatial organization, where various compartments harbor condensates that form by phase separation. Cells can control the emergence of these condensates by affecting compartment size, the amount of the involved molecules, and their physical interactions. While physical interactions directly affect compartment binding and phase separation, they can also cause oligomerization, which has been suggested as a control mechanism. Analyzing an equilibrium model, we illustrate that oligomerization amplifies compartment binding and phase separation, which reinforce each other. This nonlinear interplay can also induce multistability, which provides additional potential for control. Our work forms the basis for deriving thermodynamically consistent kinetic models to understand how biological cells can regulate phase separation in their compartments.
@@ I Introduction
Biological cells face the difficult challenge of controlling the spatial organization of their biomolecules. They partly meet this challenge by compartmentalization, where lipid membranes or other well-defined structures physically separate the cytosol into chemically-distinct regions. Within each compartment, spontaneous accumulation of biomolecules, in particular via phase separation [1], can impart further organization. Examples for these hierarchical organizations encompass compartments of various spatial dimension; see Fig. 1a. One example for a linear, 1D compartment is the synaptonemal complex (SC), a liquid-crystalline compartment forming between homologous chromosomes during meiosis [2], which binds molecules that can then diffuse along it and accumulate in small foci [3, 4, 5]. Protein accumulation on DNA might be another example for a similar 1D structure [6]. Less speculative is the formation of protein-rich domains in 2D lipid membranes [7]. These domains are thought to form by phase separation [8, 9, 10, 11], and the involved proteins can exchange with the bulk cytosol [12] or with separated membrane regions [13, 14, 15]. Finally, there are many examples of 3D compartments, either membrane-bound, like the nucleus, or membrane-less, like the nucleolus, which can harbor phase-separating proteins [16]. These examples show that cells employ an hierarchical organization, combining spontaneous phase separation with binding to existing compartments.
@@
Biological cells also face the difficulty to respond to external stimuli and deploy appropriate responses. In our context, this implies that cells control phase separation in compartments. While they can regulate the amount of material to control whether phase separation takes place [1], a faster, and presumably more efficient, approach is to regulate the physical properties of the involved molecules, e.g., by post-translational modifications [17]. On the one hand, molecular properties directly affect the affinity toward a compartment and phase separation. On the other hand, specific physical interactions between proteins might lead to the formation of small oligomers, which could affect these processes. Indeed, there is experimental evidence that oligomerization regulates protein domains on membranes [18, 19, 20, 9]. The aim of this paper is to disentangle the respective contributions of compartment binding, unspecific interactions, and oligomerization on phase separation in compartments.
@@
Biological cells are alive, implying that many processes are driven away from equilibrium, which could also affect protein domains in compartments. Previous work has described such systems using kinetic models [21, 22], reaction-diffusion equations [23, 24, 25, 26], and coupled Cahn-Hilliard equations [27, 28]. However, these phenomenological models do not take into account thermodynamic constraints, so they cannot reveal the precise influence of the physical properties of the involved molecules. Thermodynamically-consistent models have been derived for the specific case of wetting and pre-wetting on a membrane with surface binding [29, 30], but a more general theory is missing. In this paper, our aim is to unveil thermodynamic principles of phase separation in compartments based on an equilibrium model; see Fig. 1b. After introducing the model, we discuss in the subsequent sections the interplay between compartment binding, unspecific interactions, and dimerization as an example of oligomerization.
@@ II Equilibrium model of phase separation in a compartment
where $N_{\mathrm{tot}}$ is the total number of solute molecules with molecular volumes $\nu_{\mathrm{c}}$ and $\nu_{\mathrm{b}}$ in the compartment and bulk, respectively. Introducing the total fraction of solute in the compartment, $\phi=\phi_{1}+\phi_{2}$, and the total fraction in the bulk, $\psi=\psi_{1}+\psi_{2}$, material conservation becomes
@@
where $\bar{\phi}=V_{\mathrm{c}}^{-1}\int\phi\,\mathrm{d}^{n}r$ and $\bar{\psi}=V_{\mathrm{b}}^{-1}\int\psi\,\mathrm{d}^{m}r$ are the average total fractions in the respective subsystems. The relative size of the subsystems is $\eta=(\nu_{\mathrm{c}}V_{\mathrm{b}})/(\nu_{\mathrm{b}}V_{\mathrm{c}})$, where $\eta>1$ when the bulk is larger than the compartment. The fixed normalized total solute amount $\Phi_{\mathrm{tot}}=N_{\text{tot}}\nu_{\mathrm{c}}/V_{\mathrm{c}}$ denotes the fraction of the compartment occupied by solute if the bulk was empty. Note that this fraction can exceed $1$ if there is more material than can fit in the compartment.
@@
where we consider the free energy densities
@@
$\quad+\omega_{1,\mathrm{b}}\psi_{1}+\omega_{2,\mathrm{b}}\psi_{2}+\chi\psi_{\mathrm{}}(1-\psi)\Bigr{]}\;,$ (4b)
which each account for translational entropy (respective first three terms), contributions of monomers (fourth term) and dimers (fifth term), as well as an unspecific interaction between all solute molecules captured by the Flory parameter $\chi$ [31, 32]. For simplicity, we assume that the unspecific interactions are the same in both subsystems. Likewise, we assume that dimerization and compartment binding are decoupled, so that the total energy decreases by $\omega_{2}$ when a solute molecule becomes part of a dimer, independent of the subsystem it is in. This also implies that the energy $\Delta\omega$ gained when a solute molecule enters the compartment is independent of whether this molecule is part of a dimer. This choice implies $\omega_{1,\mathrm{b}}=0$, $\omega_{2,\mathrm{b}}=-\omega_{2}$, $\omega_{1,\mathrm{c}}=-\Delta\omega$, and $\omega_{2,\mathrm{c}}=-\omega_{2}-\Delta\omega$. Taken together, there are five independent parameters: $\eta$ determines the relative sizes of the compartment, $\Phi_{\mathrm{tot}}$ sets the total amount of material, and the three interaction energies $\omega_{2}$, $\Delta\omega$, and $\chi$ govern the tendency of the solute to dimerize, be located in the compartment, and phase separate, respectively; see Fig. 1b.
@@ III.1 Dimerization regulates compartment binding
We first study the interplay between dimerization and compartment binding, neglecting unspecific interactions ($\chi=0$). In this case, phase separation cannot take place and both subsystems remain homogeneous, so the free energy $F$ given by Eq. (3) only depends on the average fractions $\bar{\phi}_{1}$, $\bar{\phi}_{2}$, $\bar{\psi}_{1}$, and $\bar{\psi}_{2}$. At equilibrium, $F$ is minimized under the material-conservation constraint given by Eq. (2). We minimize $F$ by first minimizing the two terms in Eq. (3) separately, which corresponds to equilibrating the dimerization reaction. We can then express the monomer fraction $\bar{\phi}_{1}$ and the dimer fraction $\bar{\phi}_{2}$ in the compartment as a function of the total fraction $\bar{\phi}$,
@@
introducing the dimerization constant $K_{\mathrm{d}}=e^{2\omega_{2}+1}$ governing the dimerization equilibrium, $\bar{\phi}_{2}/\bar{\phi}_{1}^{2}=K_{\mathrm{d}}$; see App. A. Eq. (5) implies that the compartment is dominated by monomers in the limit $K_{\mathrm{d}}\bar{\phi}\ll 1$, whereas it contains mostly dimers when $K_{\mathrm{d}}\bar{\phi}\gg 1$. Moreover, $\bar{\phi}_{2}$ increases faster than $\bar{\phi}_{1}$ as a function of the total fraction $\bar{\phi}$, implying that the ratio of dimers increases with $\bar{\phi}$; see Fig. 2a. Analogous expressions apply for the fractions $\bar{\psi}_{1}$, $\bar{\psi}_{2}$, and $\bar{\psi}$ in the bulk. For both subsystems, we can then express the free energy density as a function of $\bar{\phi}$ and $\bar{\psi}$ alone. Since these two quantities are linked by the constraint given by Eq. (2), we can express $F$ as a function of only the bulk fraction $\bar{\psi}$. Minimizing $F$ with respect to $\bar{\psi}$ then yields the global equilibrium, where the dimerization in each subsystem and compartment binding are equilibrated.
@@
We start by discussing the limiting cases of mostly monomers ($\omega_{2}\rightarrow-\infty$). In this case, we have $\bar{\phi}_{1}=\bar{\phi}$ and $\bar{\psi}_{1}=\bar{\psi}$, whereas $\bar{\phi}_{2}=\bar{\psi}_{2}=0$. If the bulk is dilute ($\bar{\psi}_{1}\ll 1$), we find $\bar{\phi}_{1}/\bar{\psi}_{1}\approx(1-\bar{\phi}_{1})e^{\Delta\omega}$. If the compartment is also dilute ($\bar{\phi}_{1}\ll 1$), this implies that the ratio $\bar{\phi}_{1}/\bar{\psi}_{1}$ is given by $e^{\Delta\omega}$; see Fig. 2b. However, if the overall amount of material increases sufficiently ($\Phi_{\mathrm{tot}}>\frac{1}{2}+\eta e^{-\Delta\omega}$), the compartment is more than half filled ($\bar{\phi}_{1}>\frac{1}{2}$), and the ratio $\bar{\phi}_{1}/\bar{\psi}_{1}$ significantly decreases below $e^{\Delta\omega}$. For much larger $\Phi_{\mathrm{tot}}$, the compartment saturates, so that excess amount is constrained to the bulk. In the opposing limit of mostly dimers ($\omega_{2}\rightarrow\infty$), we find $\bar{\phi}_{1}=\bar{\psi}_{1}=0$, and $\bar{\phi}_{2}/\bar{\psi}_{2}=(1-\bar{\phi}_{2})^{2}e^{2\Delta\omega}$, implying an increased affinity toward the compartment compared to the monomeric case ($\bar{\phi}_{2}/\bar{\psi}_{2}>\bar{\phi}_{1}/\bar{\psi}_{1}$); see Fig. 2b. Note that this increased affinity can be interpreted in two complementary pictures: Focusing on individual particles, dimerization reduces their translational entropy, and thus the overall free energy, implying that dimers are favored. In contrast, if we focus on the number concentrations of monomers and dimers, we simply find that the binding energy of dimers is twice that of monomers, explaining their increased affinity. Taken together, binding and dimerization increase the fractions $\phi_{i}$ in the compartment until saturation is reached.
@@
We next focus on the intermediate case, where monomers and dimers can both be present in significant amounts. For instance, in the case of $\omega_{2}=3$, Fig. 2b shows that the total fractions $\bar{\phi}$ and $\bar{\psi}$ in the two subsystems transition from the monomeric limit to the dimeric limit when the total fraction $\Phi_{\mathrm{tot}}$ is increased. This transition is a consequence of the preferred binding toward the compartment and the non-linear response of dimerization: If the total fraction $\Phi_{\mathrm{tot}}$ is low, both subsystems contain mostly monomers; see Eq. (5) and Fig. 2a. Since the fraction in the compartment is larger than the fraction in the bulk ($\bar{\phi}>\bar{\psi}$), the compartment contains more dimers than the bulk, and will thus eventually transition into a state dominated by dimers while the bulk is still dominated by monomers. Finally, for sufficiently large $\Phi_{\mathrm{tot}}$, the bulk will also contain mostly dimers and the transition is completed. These three phases are summarized in Fig. 2c, indicating that we can predict the transition points by appropriate equilibration of the limiting cases; see App. B. Note that the saturation of the compartment implies another non-linearity for very large $\Phi_{\mathrm{tot}}$. Taken together, the interplay between dimerization and compartment binding leads to non-linear behavior, suggesting that dimerization, as well as the relative size $\eta$, can be used to control compartment binding (see Fig. 2d).
@@
Figure 2: Dimerization regulates compartment binding. (a) Fraction $\bar{\phi}_{2}$ of dimers normalized to total fraction $\bar{\phi}$ as as a function of $\bar{\phi}$ for a single compartment for various dimerization energies $\omega_{2}$. (b) $\bar{\phi}$ in the compartment as a function of the total fraction $\bar{\psi}$ in the bulk for $\omega_{2}=3$. The predictions for systems exclusively comprising monomers and dimers are shown as red dashed lines. (c) $\bar{\phi}$ as a function of the total fraction $\Phi_{\mathrm{tot}}$ in the system for $\omega_{2}=3$. The three scaling regimes discussed in the main text are highlighted by black lines. (d) $\bar{\phi}$ as a function of $\omega_{2}$ for various $\eta$ at $\Phi_{\mathrm{tot}}=1.5$. (a)–(d) Additional model parameters are $\eta=100$ and $\Delta\omega=2$.

### v2

@@ Abstract
Biological cells exhibit a hierarchical spatial organization, where various compartments harbor condensates that form by phase separation. Cells can control the emergence of these condensates by affecting compartment size, the amount of the involved molecules, and their physical interactions. While physical interactions directly affect compartment binding and phase separation, they can also cause oligomerization, which has been suggested as a control mechanism. Analyzing an equilibrium model, we illustrate that dimerization amplifies compartment binding and phase separation, which reinforce each other. This nonlinear interplay can also induce multistability, which provides additional potential for control. Our work forms the basis for deriving thermodynamically consistent kinetic models to understand how biological cells can regulate phase separation in their compartments.
@@ I Introduction
Biological cells face the difficult challenge of controlling the spatial organization of their biomolecules. They partly meet this challenge by compartmentalization, where lipid membranes or other well-defined structures physically separate the cytosol into chemically-distinct regions. Within each compartment, spontaneous accumulation of biomolecules, in particular via phase separation [1, 2], can impart further organization. Examples for these hierarchical organizations encompass compartments of various spatial dimension; see Fig. 1a. One example for a linear, 1D compartment is the synaptonemal complex (SC), a liquid-crystalline compartment forming between homologous chromosomes during meiosis [3], which binds molecules that can then diffuse along it and accumulate in small foci [4, 5, 6]. Protein accumulation on DNA might be another example for a similar 1D structure [7]. Less speculative is the formation of protein-rich domains in 2D lipid membranes [8]. These domains are thought to form by phase separation [9, 10, 11, 12], and the involved proteins can exchange with the bulk cytosol [13] or with separated membrane regions [14, 15, 16]. Finally, there are many examples of 3D compartments, either membrane-bound, like the nucleus, or membrane-less, like the nucleolus, which can harbor phase-separating proteins [17]. These examples show that cells employ an hierarchical organization, combining spontaneous phase separation with binding to existing compartments.
@@
Biological cells also face the difficulty to respond to external stimuli and deploy appropriate responses. In our context, this implies that cells control phase separation in compartments. While they could regulate the amount of material to control whether phase separation takes place [2, 1], a faster, and presumably more efficient, approach is to regulate the physical properties of the involved molecules, e.g., by post-translational modifications [18]. On the one hand, molecular properties directly affect the affinity toward a compartment and phase separation. On the other hand, specific physical interactions between proteins might lead to the formation of small oligomers, which could affect these processes. Indeed, there is experimental evidence that oligomerization regulates protein domains on membranes [19, 20, 21, 11]. The aim of this paper is to disentangle the respective contributions of compartment binding, unspecific interactions, and oligomerization on phase separation in compartments.
@@
Biological cells are alive, implying that many processes are driven away from equilibrium, which could also affect protein domains in compartments. Previous work has described such systems using kinetic models [22, 23], reaction-diffusion equations [24, 25, 26, 27], and coupled Cahn-Hilliard equations [28, 29]. However, these phenomenological models do not take into account thermodynamic constraints, so they cannot reveal the precise influence of the physical properties of the involved molecules. Thermodynamically consistent models have been derived for the specific case of wetting and pre-wetting on a membrane with surface binding [30, 31], but a more general theory is missing. As a step toward such a general theory, we here focus on the behavior within the compartment by discussing an equilibrium theory that captures the interplay between compartment binding, unspecific interactions enabling phase separation, and dimerization as an example of oligomerization; see Fig. 1b.
@@ II Equilibrium model of phase separation in a compartment
where $N_{\mathrm{tot}}$ is the total number of solute molecules with molecular volumes $\nu_{\mathrm{c}}$ and $\nu_{\mathrm{b}}$ in the compartment and bulk, respectively. Introducing the fraction of solute in the compartment, $\phi=\phi_{1}+\phi_{2}$, and the solute fraction in the bulk, $\psi=\psi_{1}+\psi_{2}$, material conservation becomes
@@
where $\bar{\phi}=V_{\mathrm{c}}^{-1}\int\phi\,\mathrm{d}^{n}r$ and $\bar{\psi}=V_{\mathrm{b}}^{-1}\int\psi\,\mathrm{d}^{m}r$ are the average solute fractions in the respective subsystems. The key parameter $\eta=(\nu_{\mathrm{c}}V_{\mathrm{b}})/(\nu_{\mathrm{b}}V_{\mathrm{c}})$ quantifies the size of the bulk relative to the compartment, where here size is measured by the maximal number of solutes that can fit in the respective subsystem. Without loss of generality, we assume that the bulk is larger, $\eta>1$, essentially defining which subsystem is the bulk. Typical values are $\eta\sim 115$ in the PAR system of C. elegans [32, 19] and $\eta\sim 90$ for the synaptonemal complex in A. thalania (private communication), so we use $\eta=100$ in this paper. The final parameter in Eq. (2) is the fixed normalized total solute amount $\Phi_{\mathrm{tot}}=N_{\text{tot}}\nu_{\mathrm{c}}/V_{\mathrm{c}}$, which denotes the fraction of the compartment occupied by solute if the bulk was empty. Note that this fraction can exceed $1$ if there is more material than can fit in the compartment.
@@
where we neglect coupling terms describing wetting since we focus on the behavior inside the compartment. Such couplings could also affect phase separation when the compartment’s dimension is lower than the bulk’s, but this typically just leads to rescaled parameters; see Eq. (19) in Appendix A. We thus focus on the bulk effects inside the subsystems, which are only coupled via material exchange. The equilibrium of this situation is governed by the respective free energy densities accounting for dimerization and phase separation,
@@
$\quad+\omega_{1,\mathrm{b}}\psi_{1}+\omega_{2,\mathrm{b}}\psi_{2}+\chi\psi_{\mathrm{}}(1-\psi)\Bigr{]}\;.$ (4b)
For each expression, the respective first three terms describe translational entropies, whereas the other terms account for enthalpic contributions of monomers (fourth term) and dimers (fifth term), as well as an unspecific interaction between all solute molecules captured by the Flory parameter $\chi$ [33, 34]. For simplicity, we assume that solute monomers and solvent molecules have the same molecular size, e.g., because they are both macromolecules. To clearly separate the influence of physical interactions and limit the number of physical parameters, we consider the case where unspecific interactions are the same in both subsystems, and that dimerization and unspecific interactions are independent. Likewise, we consider the case where dimerization and compartment binding are decoupled, so that the total energy decreases by $\omega_{2}$ when a solute molecule becomes part of a dimer, independent of the subsystem it is in. This also implies that the energy $\Delta\omega$ gained when a solute molecule enters the compartment is independent of whether this molecule is part of a dimer. These choices imply $\omega_{1,\mathrm{b}}=0$, $\omega_{2,\mathrm{b}}=-\omega_{2}$, $\omega_{1,\mathrm{c}}=-\Delta\omega$, and $\omega_{2,\mathrm{c}}=-\omega_{2}-\Delta\omega$.
@@
The behavior of the system is governed by the balance of chemical potentials derived from Eqs. (4) together with the constraint given by Eq. (2). This implies that neither the dimensions of the subsystems nor their absolute volumes affect the behavior, which is instead governed by the following five parameters: $\eta$ determines the relative sizes of the subsystems, $\Phi_{\mathrm{tot}}$ sets the total amount of material, and the three interaction energies $\omega_{2}$, $\Delta\omega$, and $\chi$ govern the tendency of the solute to dimerize, be located in the compartment, and phase separate, respectively; see Fig. 1b.
@@ III.1 Dimerization regulates compartment binding
We first study the interplay between dimerization and compartment binding, neglecting unspecific interactions ($\chi=0$). In this case, phase separation cannot take place and both subsystems remain homogeneous, so the free energy $F$ given by Eq. (3) only depends on the average fractions $\bar{\phi}_{1}$, $\bar{\phi}_{2}$, $\bar{\psi}_{1}$, and $\bar{\psi}_{2}$. To minimize $F$ under the material-conservation constraint given by Eq. (2), we first minimize the two terms in Eq. (3) separately, which corresponds to equilibrating the dimerization reaction. We can then express the monomer fraction $\bar{\phi}_{1}$ and the dimer fraction $\bar{\phi}_{2}$ in the compartment as a function of the solute compartment fraction $\bar{\phi}$,
@@
introducing the dimerization constant $K_{\mathrm{d}}=e^{2\omega_{2}+1}$ governing the dimerization equilibrium, $\bar{\phi}_{2}/\bar{\phi}_{1}^{2}=K_{\mathrm{d}}$; see Appendix B. Eq. (5) implies that the compartment is dominated by monomers in the limit $K_{\mathrm{d}}\bar{\phi}\ll 1$, whereas it contains mostly dimers when $K_{\mathrm{d}}\bar{\phi}\gg 1$. Moreover, $\bar{\phi}_{2}$ increases faster than $\bar{\phi}_{1}$ as a function of the solute compartment fraction $\bar{\phi}$, implying that the ratio of dimers increases with $\bar{\phi}$; see Fig. 2a. Analogous expressions apply for the fractions $\bar{\psi}_{1}$, $\bar{\psi}_{2}$, and $\bar{\psi}$ in the bulk. For both subsystems, we can then express the free energy density as a function of $\bar{\phi}$ and $\bar{\psi}$ alone. Since these two quantities are linked by the constraint given by Eq. (2), we can express $F$ as a function of only the solute bulk fraction $\bar{\psi}$. Minimizing $F$ with respect to $\bar{\psi}$ then yields the global equilibrium, where the dimerization in each subsystem and compartment binding are equilibrated.
@@
We start by discussing the limiting cases of mostly monomers ($\omega_{2}\rightarrow-\infty$). In this case, we have $\bar{\phi}_{1}=\bar{\phi}$ and $\bar{\psi}_{1}=\bar{\psi}$, whereas $\bar{\phi}_{2}=\bar{\psi}_{2}=0$. If the bulk is dilute ($\bar{\psi}_{1}\ll 1$), we find $\bar{\phi}_{1}/\bar{\psi}_{1}\approx(1-\bar{\phi}_{1})e^{\Delta\omega}$. If the compartment is also dilute ($\bar{\phi}_{1}\ll 1$), this implies that the ratio $\bar{\phi}_{1}/\bar{\psi}_{1}$ is given by $e^{\Delta\omega}$; see Fig. 2b. However, if the overall amount of material increases sufficiently ($\Phi_{\mathrm{tot}}>\frac{1}{2}+\eta e^{-\Delta\omega}$), the compartment is more than half filled ($\bar{\phi}_{1}>\frac{1}{2}$), and the ratio $\bar{\phi}_{1}/\bar{\psi}_{1}$ significantly decreases below $e^{\Delta\omega}$. For much larger $\Phi_{\mathrm{tot}}$, the compartment saturates, so that excess amount is constrained to the bulk. In the opposing limit of mostly dimers ($\omega_{2}\rightarrow\infty$), we find $\bar{\phi}_{1}=\bar{\psi}_{1}=0$, and $\bar{\phi}_{2}/\bar{\psi}_{2}=(1-\bar{\phi}_{2})^{2}e^{2\Delta\omega}$, implying an increased affinity toward the compartment compared to the monomeric case ($\bar{\phi}_{2}/\bar{\psi}_{2}>\bar{\phi}_{1}/\bar{\psi}_{1}$); see Fig. 2b. Note that this increased affinity can be interpreted in two complementary pictures: Focusing on individual particles, dimerization reduces their translational entropy, thus requiring less binding energy to concentrate them in the compartment. In contrast, if we focus on the number concentrations of monomers and dimers, we simply find that the binding energy of dimers is twice that of monomers, explaining their increased affinity. Taken together, binding and dimerization increase the fractions $\phi_{i}$ in the compartment until saturation is reached.
@@
We next focus on the intermediate case, where monomers and dimers can both be present in significant amounts. For instance, in the case of $\omega_{2}=3$, Fig. 2b shows that the solute fractions $\bar{\phi}$ and $\bar{\psi}$ in the two subsystems transition from the monomeric limit to the dimeric limit when the total fraction $\Phi_{\mathrm{tot}}$ is increased. This transition is a consequence of the preferred binding toward the compartment and the non-linear response of dimerization: If the total fraction $\Phi_{\mathrm{tot}}$ is low, both subsystems contain mostly monomers; see Eq. (5) and Fig. 2a. Since the fraction in the compartment is larger than the fraction in the bulk ($\bar{\phi}>\bar{\psi}$), the compartment contains more dimers than the bulk, and will thus eventually transition into a state dominated by dimers while the bulk is still dominated by monomers. Finally, for sufficiently large $\Phi_{\mathrm{tot}}$, the bulk will also contain mostly dimers and the transition is completed. These three phases are summarized in Fig. 2c, indicating that we can predict the transition points by appropriate equilibration of the limiting cases; see Appendix C. Note that the saturation of the compartment implies another non-linearity for very large $\Phi_{\mathrm{tot}}$. Taken together, the interplay between dimerization and compartment binding leads to non-linear behavior, which has been observed experimentally [19]. This behavior suggests that dimerization, as well as the relative size $\eta$, can be used to control compartment binding.
@@
Figure 2: Dimerization regulates compartment binding. (a) Fraction $\bar{\phi}_{2}$ of dimers normalized to solute compartment fraction $\bar{\phi}$ as as a function of $\bar{\phi}$ for a single compartment for various dimerization energies $\omega_{2}$. (b) $\bar{\phi}$ in the compartment as a function of the solute fraction $\bar{\psi}$ in the bulk for $\omega_{2}=3$. The predictions for systems exclusively comprising monomers and dimers are shown as red dashed lines. (c) $\bar{\phi}$ as a function of the total fraction $\Phi_{\mathrm{tot}}$ in the system for $\omega_{2}=3$. The three scaling regimes discussed in the main text are highlighted by black lines. (d) $\bar{\phi}$ as a function of $\omega_{2}$ for various $\eta$ at $\Phi_{\mathrm{tot}}=1.5$. (a)–(d) Additional model parameters are $\eta=100$ and $\Delta\omega=2$.